OpenStudy (anonymous):

Equation for an ellipse with center (0,0), vertices (3,0) and (-3,0), and foci (2,0) and (-2,0)?

OpenStudy (anonymous):

a. x^2/9+y^2/4=1 b. x^2/5+y^2/9=1 c. x^2/9+y^2/5=1 d. x^2/4+y^2/9=1

OpenStudy (nottim):

dunno what a foci is

OpenStudy (nottim):

really, i dunno what this is in general

OpenStudy (nottim):

may i offer the services of wolfram alpha to help?

OpenStudy (anonymous):

At this point PLEASE!

OpenStudy (anonymous):

Can you provide me with a link

OpenStudy (amistre64):



OpenStudy (amistre64):

center to vertex = 3 center to focus = 2 a^2 = 3^2 + 2^2 = 13 a = sqrt(13); so we should have a 9 and a 13 in the denoms

OpenStudy (amistre64):

hmmm

OpenStudy (amistre64):

the denoms measure the 1/2 axis after we sqrt them

OpenStudy (amistre64):

sqrt(9) = 3 sqrt(n) = sqrt(13); n = 13 what am i missing?

OpenStudy (anonymous):

Hmm... I'm just confused. But I understand your useful explanation

OpenStudy (amistre64):

 i had my verts in the wrong spots :)

OpenStudy (amistre64):

9-4 = 5 sqrt(5)^2 = 5 x/9 and y/5 should do it

OpenStudy (anonymous):

Oh o.k. Now it makes sense. Thanks amistre64. You're the best

OpenStudy (amistre64):

good luck